Let the production function of a firm is given as

π‘ž=(π‘₯0.5 +𝑦0.5)2

Where π‘₯ and 𝑦 are inputs and 𝑀π‘₯ is the price of input π‘₯ and 𝑀𝑦 is the price

of input 𝑦.

a) Assume the firm has a limited budget to spend on buying input. Find

the cost-conditional input demand function for each input.

b) Find the cost function of the firm


1
Expert's answer
2021-10-12T13:47:39-0400

Solution:

a.). Derive MRTS =Β "\frac{MP_{X} }{MP_{Y}}"


MPXΒ =Β "\frac{\partial Q} {\partial X} = \frac{1} {x^{0.5} }"


MPYΒ =Β "\frac{\partial Q} {\partial Y} = \frac{1} {y^{0.5} }"


"\frac{MP_{X} }{MP_{Y}} = \frac{w }{r}"


"\frac{\frac{1} {x^{0.5} } }{\frac{1} {y^{0.5} } } = \frac{w} {r }"


"\frac{y^{0.5}} {x^{0.5} } = \frac{w} {r }"


x =Β "\frac{r^{2}y} {w^{2} }"


y =Β "\frac{w^{2}x} {r^{2} }"

Β 

b.). The cost function of the firm:

TC = rY + wX

TC = r("\frac{w^{2}x} {r^{2} }") + w("\frac{r^{2}y} {w^{2} }")

TC =Β "\frac{w^{2}x} {r} +"Β "\frac{r^{2}y} {w}"